OpenStudy (anonymous):

A ball is thrown straight into the air from 80 feet above ground level with an initial velocity of 40 feet per second. Give answers correct to the nearest hundredth. When is the height of the ball 100 feet? Explain how you used your graphing calculator to find the answer.

OpenStudy (michele_laino):

please, are you asking for the final velocity?

OpenStudy (michele_laino):

hint: since the gravitational field is a conservative field, then total energy (kinetic energy plus potential energy) has to be constant, so we can write: \[\Large \frac{1}{2}mv_f^2 + mg{h_f} = \frac{1}{2}mv_i^2 + mg{h_i}\] or: \[\Large \frac{1}{2}v_f^2 + g{h_f} = \frac{1}{2}v_i^2 + g{h_i}\] where \(v_i=40\), \(h_i=80\), \(h_f=100\), \(v_f\) is the magnitude of the final velocity, and \(g=32\) \(feet/sec^2\)